OpenStudy (anonymous):

When 3n is divided by 7, the remainder is 4. What is the remainder when 2n is divided by 7? Help, a bit confused here...again. Please guys

OpenStudy (nirmalnema):

when 3n is divided by 7 its not always gives 4 eg.. n = 3 or 4 its gives 2 and 5

OpenStudy (anonymous):

@nirmalnema wait, so, ya, if 3n is divided by 7 and the remainder is 4, so that means n = 6?

OpenStudy (nirmalnema):

i dont said that it doesn't give .. i said it not always correct

OpenStudy (the_fizicx99):

Yes, n would = 6 As 6*3 = 18 7 goes into 18 2 times and the remainder is 4

OpenStudy (anonymous):

@nirmalnema ahh ok, but i got a bit of your point. I tried n = 6 (trial&error) and did all the equation stuff and I got 5 as the remainder. I have answer key here though

OpenStudy (the_fizicx99):

How did you get 5 lol ._.

OpenStudy (anonymous):

@tHe_FiZiCx99 the question asked for the remainder. Since the divided is 2n and the divisor is 7, ---- 2(6) divided by 7 is equal to 1 r 5. =))

OpenStudy (nirmalnema):

yup its correct...

OpenStudy (anonymous):

ah ok btw thanks for your point :)

OpenStudy (nirmalnema):

your welcome... = P

OpenStudy (the_fizicx99):

Oh I was taking about 3n divided by 7 xD

ganeshie8 (ganeshie8):

you may lessen the effort during trial/error if u check both positive and negative remainders :- 3n = 4 (mod 7) 3(-1) = 4 (mod 7) n = -1 (mod 7) => 2n = -2 (mod 7) = 5 (mod 7)